OpenStudy (anonymous):

Solve for x:(x+4)/3 = 2

OpenStudy (anonymous):

\[{(x+4)\over3}=2\]get rid of the \(3\) first by multiplying both sides by \(3\):\[{\cancel3\over1}\times{(x+4)\over\cancel3}=2\times3\]\[x+4=6\]\[x=\underline{~~~}?\]

OpenStudy (anonymous):

x=2?

OpenStudy (anonymous):

yup :)